# SWAN: Preprocessing SGD Enables Adam-Level Performance On LLM Training With Significant Memory Reduction

### v1

@@ 3.1 SWAN Update Rules
In SWAN (Algorithm 1), the raw SGD gradient ${\mathbf{G}}^{t}$ is processed by the following operations  [^1] Here ${\mathbf{G}}$ is assumed to be the gradient matrix for some parameter block in the model (e.g. a linear layer); enabling us to take advantage of the matrix structure to achieve our smoothing and whitening goals.:
@@
$\mathtt{GradNorm}$: gradient invariance & stabilization.
@@
Consider the gradient matrix ${\mathbf{G}}\in\mathbb{R}^{m\times n}$ with rows and columns corresponding to input and output dimensions, respectively, of some block of model parameters. Let $1\leq i\leq m$ represent the input indices and $1\leq j\leq n$ represent the output indices. Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, we propose to standardization across the output dimensions at each time step $t$, defined as:
$\mathtt{GradNorm}({\mathbf{G}})=\frac{{\mathbf{G}}-\bar{{\bm{g}}}\mathbf{1}_{n}^{\top}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
@@
where $\bar{{\bm{g}}}=\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}$ is the mean across output dimensions at time $t$; ${\bm{s}}=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\bar{{\bm{g}}})^{2}}$ is the standard deviation across dimensions at time $t$; $\bar{{\bm{g}}}$ and ${\bm{s}}$ are ${m}$-dimensional column vectors; and $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
@@
$\mathtt{GradNorm}$ is an example of operators typically used in the forward pass applied to the backward pass: $\mathtt{GradNorm}$ is LayerNorm (Ba et al., 2016) (without learnable parameters) being applied to the backward gradient rather than forward activations. Similar to Layer Norm, $\mathtt{GradNorm}$ allows the optimizer to be invariant under the following transformations on gradient matrix: i) matrix-wise re-scaling; ii) matrix-wise re-centering; iii) row-wise re-scaling; and iv) row-wise re-centering. Compared to the sign operation ($\text{sign}({\mathbf{G}})$) and matrix-wise normalization ($\frac{{\mathbf{G}}}{\|{\mathbf{G}}\|}$) used in signed descent, $\mathtt{GradNorm}$ preserves richer information of the gradient scaling while offering invariance properties. Moreover, in Section 4 we will show that when applied to a transformer architecture, $\mathtt{GradNorm}$ also helps stabilize the gradient distribution across time.
$\mathtt{GradWhitening}$: gradient whitening.
Adam performs approximate gradient whitening element-wise, leading to well-conditioned regions. In contrast, we propose the matrix $\mathtt{GradWhitening}$ operator, which performs whitening at the matrix level. Given the gradient ${\mathbf{G}}$, the $\mathtt{GradWhitening}$ operator simply orthogonalizes ${\mathbf{G}}$  [^2] An equivalent definition is given by $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{U}}{\mathbf{V}}^{\top}$, where ${\mathbf{U}}$ and ${\mathbf{V}}$ are both orthogonal matrices defined via SVD decomposition of ${\mathbf{G}}={\mathbf{U}}{\mathbf{S}}{\mathbf{V}}^{\top}$. :
$\mathtt{GradWhitening}({\mathbf{G}}):=({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ (2)
@@
where the exponent $-\frac{1}{2}$ stands for matrix inverse square root. One can show that $({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ is simply the orthogonalization of ${\mathbf{G}}$, i.e., the closest orthogonal matrix to ${\mathbf{G}}$ (w.r.t the Frobenius norm).
@@ 3.2 Practical Considerations
Various GPU-friendly and fast approximations are developed in the literature of decorrelated batch normalization or feature whitening. We consider the naive Netwon-Schulz (N-S) iteration algorithm (Li et al., 2018; Huang et al., 2019) to compute $\mathtt{GradWhitening}$:
@@
[…]
@@
The whitening operation $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, hence might drastically change the effective learning rate. In practice, we propose the following re-scaling before updating the weight:
@@
This will helps us to rescale the norm of whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, and allowing us to get reasonably good performance without extensively tuning learning rates–we found that the learning rates used Adam usually also work well for SWAN. It is also possible to remove rescaling, but that would require additional hyper-parameter search to find the optimal learning-rate for SWAN.
@@ 4 Analysis: A LLM Learning Dynamics Perspective
As a new stateless adaptive optimizer, the complete theoretical properties of SWAN is an open question which we leave for future work. However, as a first analysis, we consider SWAN from a learning dynamics perspective, specifically the dynamics of an LLM based upon a simplified transformer block. It is this analysis that led to the design of SWAN.
@@ 4.1 Setup
We assume the following simplified transformer block (STB) architecture recently proposed in Tian et al. (2023):
@@
4.2 $\mathtt{GradNorm}$ stabilizes gradient distributions of simplified transformers
[…]
@@ 4.2 $\mathtt{GradNorm}$ stabilizes gradient distributions of simplified transformers
4.3 $\mathtt{GradWhitening}$ Is An Efficient Non-diagonal Second-Order Update
In this section, we show that $\mathtt{GradWhitening}$ is equivalent to non-diagonal approximate second-order method, under a specific Kronecker factorization assumption of the Hessian. The assumption is as below:
@@ 4.3 $\mathtt{GradWhitening}$ Is An Efficient Non-diagonal Second-Order Update
By leveraging this structural assumption, we can now effectively estimate ${\mathbf{H}}$ by only using one single gradient matrix sample $\bm{G}:=[{\bm{g}}_{1},\dots,{\bm{g}}_{n}]\in\mathbb{R}^{m\times n}$. Recall that the Fisher information formulation of Hessian is defined as ${\mathbf{H}}=\mathbb{E}(\text{Vect}({\mathbf{G}})\text{Vect}({\mathbf{G}})^{\top})$ where $\text{Vect}(\cdot)$ denotes the vectorized operation. Under assumption 1, we can estimate ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$, using a single sample of the gradient, that is we can compute the following simple estimate $\frac{1}{n}\sum_{i=1}^{n}{\bm{g}}_{i}{\bm{g}}_{i}^{T}=\bm{G}\bm{G}^{\top}$, which approximate the current $\tilde{{\mathbf{H}}}$. Hence, $\mathtt{GradWhitening}$ can be seen as applying a second order update under our structural assumption:
@@
In Theorem 2 of Appendix, we will further show that the pre-conditioned update rule $\tilde{{\mathbf{H}}}^{-\frac{1}{2}}{\mathbf{G}}$ indeed induces a Newton-like behavior. More importantly, in the following Proposition, we show that the implicit assumption of $\mathtt{GradWhitening}$ is a good abstraction of LLM dynamics since it in fact aligns with the equilibrium Hessian structure in the STB regime:
@@
Proposition 1 shows that, under a simplified setting of the transformer, the Hessian will also converge to an equilibrium solution where the $M_{C}\times M_{C}$ blocks over the diagonal direction of Hessian shares the identical structure, which supports the implicit assumption of $\mathtt{GradWhitening}$. This result is verified in our numerical experiment (Appendix, Figure 9).
Finally, Theorem 3 presented in Appendix will also how $\mathtt{GradWhitening}$ helps to make the convergence rate of SGD more robust to the condition number of local curvatures, and outperforms both SGD and Adam in the ill-conditioned regime.
@@ 5 Experiments
In this section, we perform empirical studies on SWAN. The main goals of those studies are as follows:
  • Verifying whether $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ indeed effectively fulfilled their design goals (stabilizing gradient noise, and modeling/counteracting local curvatures). This is performed in Section 5.1 and Section 5.2.
  • Verifying whether SWAN can consistently offer Adam-level or even better performance, while maintaining low memory cost and high-throughput. This is performed in Section 5.3, Section 5.4 and Section 5.5, respectively.
  • Ablation study (Section 5.6).
@@
5.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
Figure 4: (a) - (c): KL Divergence of gradient distributions across training Steps. We use the projection weights in attention and MLP modules of the second layer as an example. The plots compare standard training with $\mathtt{GradNorm}$-augmented training. Lower KL divergence values indicate greater stability in gradient distributions. (d): Training loss curve comparing standard training vs $\mathtt{GradNorm}$-augmented training
To investigate whether $\mathtt{GradNorm}$ indeed stabilizes stochastic gradient distributions as claimed in Theorem 1, we conduct a series of controlled experiments using a scaled-down version of the LLaMA architecture. We aim to assess the impact of $\mathtt{GradNorm}$ on the gradient distribution during training by analyzing the statistical properties of gradients across multiple training steps. Specifically, we employ a small-scale LLaMA-based model with approximately 10 million parameters  (Lialin et al., 2023). Training is conducted on the C4 dataset (Raffel et al., 2020).
Baselines
We compare the following methods:
  • Standard training: this baseline utilize SGD optimizer with a learning rate of $5\times e^{-4}$, and apply a linear learning rate scheduler with a warm-up period comprising 10% of the total training steps (10,000). ]
  • $\mathtt{GradNorm}$-augmented training: apply $\mathtt{GradNorm}$ on noisy gradients before SGD update. We use the same hyperparameters as the SGD baseline.
Methodology
For each method, we measure the change of gradient statistics due to mini-batch sampling noise. For each training step, after the parameter update, we additionally sample 16 different mini-batches (each with batch size 64), and calculate the mean and standard deviation of the corresponding gradients on each of those batches. Then, we estimate the Kullback-Leibler divergence between the current gradient distribution (raw gradient and with $\mathtt{GradNorm}$) relative to the initial distribution at $t=0$. This allows us to quantify the temporal evolution of gradient distributions.
Results
Figure 4 illustrates the KL divergence of gradient distributions over training steps for both standard training and $\mathtt{GradNorm}$-augmented training, relative to the initial distribution at $t=0$. Apart from minor spikes in the beginning of training, $\mathtt{GradNorm}$ consistently and effectively stabilizes the gradient noise distribution. Furthermore, as a by-product, the loss curves demonstrate that models trained with $\mathtt{GradNorm}$ performs indeed improves upon standard training.
Ablation
One question that we might ask is this: is the performance boost of $\mathtt{GradNorm}$ simply due to that fact that $\mathtt{GradNorm}$ might be increasing the effective learning rate. We conduct detailed ablation experiments in Section 5.6, in a larger scale setting, which show this is not the case. These findings validate the hypothesis that $\mathtt{GradNorm}$ serves as an effective mechanism for mitigating the adverse effects of mini-batch noise on gradient distributions, thereby enhancing the stability and performance of the training process.
5.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
We aim to evaluate the pure optimization capability of $\mathtt{GradWhitening}$ processed gradient descent on classic optimization problems. We consider the following three settings:
(a) Quadratic optimization, well-conditioned
(b) Quadratic optimization, ill-conditioned
(c) Rastrigin function optimization
(d) Ablation on initialization
  • High-dimensional quadratic optimization. We construct quadratic optimization problem of the form in Equation 23, with ${\mathbf{W}}\in\mathbb{R}^{50\times 50}$.
  • Ill-conditioned quadratic optimization. We construct a quadratic optimization problem with the same dimensionality and formulation as above, but this time we deliberately choose an ill-conditioned ${\mathbf{H}}$.
  • Non-convex optimization with multiple local optimas. We consider the following multi-variate Rastrigin function:
    $f({\mathbf{W}})=m^{2}{\mathbf{A}}+\frac{1}{2}\text{Tr}[{\mathbf{W}}^{\top}{\mathbf{W}}]-{\mathbf{A}}\sum_{ij}\text{cos}(2\pi W_{ij})$
    where ${\mathbf{W}}$ is $m\times m$ matrix. This function has $10^{m^{2}}$ possible local optima. This is a complicated stress test for optimization algorithms, where many methods are expected to fail at different regimes. We set $m=50$, and compare the same methods as in the previous experiment.
Baselines
On all three settings (quadratic, quadratic ill-conditioned, and Rastrigin) we compare 5 different methods, including 3 baselines and 2 $\mathtt{GradWhitening}$ based variants. For baselines, we have: gradient descent (with theoretical optimal learning rate in Theorem 4); Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$ with hand-tuned learning rate); and Newton’s method with tuned learning rate. For 2 $\mathtt{GradWhitening}$-based methods, we consider: $\mathtt{GradWhitening}$-augmented GD with orthogonal initialization (i.e., ${{\mathbf{W}}^{(0)}}^{\top}{\mathbf{W}}^{(0)}={\mathbf{I}}$); and without orthogonal initialization. This is to verify whether the claims of Theorem 2 and Theorem 3 is true (i.e., $\mathtt{GradWhitening}$-processed GD behaves like Newton’s method under orthogonal initialization). All methods are initialized with the same initial condition (apart from $\mathtt{GradWhitening}$ with orthogonal initialization, which will project the original initialization onto the orthogonal version).
Results
Results are shown in Figure 5, (a) - (c). Main findings are summarized as follows:
  • On quadratic optimization problems (Figure 5 (a) and (b)), $\mathtt{GradWhitening}$ with orthogonal initialization and Newton’s method converges to optimal solution with 1 step of optimization, confirming our theoretical prediction in Theorem 2 and Theorem 3.
  • On quadratic optimization problems (Figure 5 (a)), when the problem is well-conditioned, standard GD outperforms both Adam and $\mathtt{GradWhitening}$ (non-orthogonal initialization). However, when the problem becomes ill conditioned (Figure 5 (b)), $\mathtt{GradWhitening}$ (non-orthogonal initialization) outperforms GD by a large margin; while GD suffers from convergence.
  • In all three settings (Figure 5, (a) - (c)), $\mathtt{GradWhitening}$ with non-orthogonal initialization consistently outperforms Adam, confirming our theoretical results in Proposition 3.
  • On Rastrigin function optimization (Figure 5 (c)), $\mathtt{GradWhitening}$ performs competitively to Newton’s method, with or without orthogonal initialization.
Ablation
Furthermore, we additionally conduct ablation study on the effect of initialization locations, as shown in fig. 5 (d). In this case, we force all methods to use the same orthogonalized initialization location for optimization. As shown by the result, $\mathtt{GradWhitening}$ still consistently outperforms all baselines, eliminating the potential hypothesis that orthogonalized initialization gave $\mathtt{GradWhitening}$ an unfair advantage.

### v2

@@ 4.1 SWAN Update Rules
In SWAN (Algorithm 1), the raw SGD gradient ${\mathbf{G}}^{t}$ is processed by the operations below [^1] Here ${\mathbf{G}}$ is assumed to be the gradient matrix for some parameter block in the model (e.g. a linear layer); enabling us to take advantage of the matrix structure to achieve our smoothing and whitening goals.:
@@
On the $\mathtt{GradNorm}$ Step.
@@
Consider the gradient matrix ${\mathbf{G}}\in\mathbb{R}^{m\times n}$ with rows and columns corresponding to input and output dimensions, respectively, of some block of model parameters. Let $1\leq i\leq m$ represent the input indices and $1\leq j\leq n$ represent the output indices. Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, as done in Adam, we propose to standardize across the output dimensions at each time step $t$, as follows:
$\mathtt{GradNorm}({\mathbf{G}}):=\frac{{\mathbf{G}}-\bar{{\bm{g}}}\mathbf{1}_{n}^{\top}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
@@
where $\bar{{\bm{g}}}:=\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}$ is the mean across output dimension; ${\bm{s}}:=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\bar{{\bm{g}}})^{2}}$ is the standard deviation across dimension; $\bar{{\bm{g}}}$ and ${\bm{s}}$ are ${m}$-dimensional column vectors; $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
@@
$\mathtt{GradNorm}$ is the forward pass operator LayerNorm (Ba et al., 2016) (without learnable parameters) applied on backward gradients. Similar operations have been considered in the literature to process gradients (You et al., 2019). $\mathtt{GradNorm}$ allows the optimizer to be invariant under the following transformations on the gradient matrix: i) matrix-wise re-scaling; ii) matrix-wise re-centering; iii) row-wise re-scaling; and iv) row-wise re-centering. Compared to the sign operation ($\text{sign}({\mathbf{G}})$) and matrix-wise normalization ($\frac{{\mathbf{G}}}{\|{\mathbf{G}}\|}$) used in signed descent, $\mathtt{GradNorm}$ preserves richer information of the gradient scaling while offering invariance properties. In Section 5 we will show that when applied to a simplified transformer architecture, $\mathtt{GradNorm}$ also helps stabilize the gradient distribution.
On the $\mathtt{GradWhitening}$ Step.
As discussed in Section 3, Adam relies on a second moment estimate to perform element-wise gradient whitening. More formally, the second moment of Adam estimate a diagonal approximation of the Fisher information matrix (FIM) via (Kingma & Ba, 2015; Hwang, 2024):
$\mathbb{E}(\text{Vec}({\mathbf{G}})\text{Vec}({\mathbf{G}})^{\top})\approx\text{Diag}[\text{Vec}[\mathbb{E}({\mathbf{G}}^{\odot 2})]]$
where $\text{Vec}(\cdot)$ denotes the vectorized operation, $\text{Diag}(\cdot)$ denotes the operation that produces a diagonal matrix given a vector, and $\odot$ denotes the Hadamard product. Hence, whitening using the inverse square root of this diagonal FIM is equivalent to element-wise rescaling with $\frac{1}{\sqrt{\mathbb{E}({\mathbf{G}}^{\odot 2})}}$. Here, we consider the following non-diagonal approximation:
$\mathbb{E}(\text{Vec}({\mathbf{G}})\text{Vec}({\mathbf{G}})^{\top})\approx{\mathbf{I}}_{n\times n}\otimes{\mathbf{G}}{\mathbf{G}}^{\top}\;,$
where $\otimes$ denotes Kronecker product. This leads to the following whitening step:
$\mathtt{GradWhitening}({\mathbf{G}}):=({\mathbf{G}}{\mathbf{G}}^{\top})^{-1/2}{\mathbf{G}}$ (2)
@@
where the exponent $-\frac{1}{2}$ stands for matrix inverse square root. $({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ is simply the orthogonalization of ${\mathbf{G}}$, i.e., the closest orthogonal matrix to ${\mathbf{G}}$ (w.r.t the Frobenius norm). The derivation of Equation 2 as structured FIM/Hessian approximation is discussed in Section 5.3.
@@ 4.2 Practical Considerations
Computing $\mathtt{GradWhitening}$ exactly can be expensive, as it involves solving the matrix square-root inverse. We hence directly apply the Newton-Schulz variant of decorrelated batch normalization (Song et al., 2022; Li et al., 2018; Huang et al., 2019), which allows a more GPU-friendly estimation. This is given by (Song et al., 2022; Li et al., 2018):
@@
where ${\mathbf{Y}}_{0}={\mathbf{G}}{\mathbf{G}}^{\top}$ , ${\mathbf{Z}}_{0}={\mathbf{I}}$. At convergence, $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{Z}}{\mathbf{G}}$ (Algorithm 2). See Appendix H for implementation details.
Note that the same N-S procedure has been discussed in (Mei et al., 2023) and  (Jackson, 2023) for preconditioned optimizers. Another N-S procedure has also been introduced in (Bernstein & Newhouse, 2024b), and optimized in (Jordan et al., 2024) under BF16 precision.
@@
The whitening operation $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, and as such might drastically change the effective learning rate. In practice, we propose the following re-scaling before updating the weight:
@@
This helps rescale the norm of the whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, and allowing us to get reasonably good performance without extensively tuning learning rates; we found that the learning rates used by Adam tend to work well for SWAN. We can remove rescaling, but doing so would require additional hyper-parameter search to find the optimal learning rate.
@@ 5 Analysis: A LLM Learning Dynamics Perspective
As a new stateless adaptive optimizer, the complete theoretical properties of SWAN is an open question which we leave for future work. However, as a first analysis, we consider SWAN from a learning dynamics perspective, specifically the dynamics of an LLM based upon a simplified transformer block. It is this analysis that led to the design choices for SWAN.
@@ 5.1 Setup
We consider the simplified transformer block (STB) architecture proposed in (Tian et al., 2023).
@@
5.2 $\mathtt{GradNorm}$: stabilizing gradient distributions of STB
Below we show that, based on the dynamics of the STB, $\mathtt{GradNorm}$ stabilizes ${\bm{\varepsilon}}_{{\mathbf{W}}}^{(t)}$.
@@ 5.2 $\mathtt{GradNorm}$: stabilizing gradient distributions of STB
5.3 $\mathtt{GradWhitening}$: An Efficient Non-diagonal Second-Order Update
Here, we show that $\mathtt{GradWhitening}$ is equivalent to a non-diagonal second-order method under a specific Kronecker factorization assumption of the Hessian/FIM. The assumption is as below:
@@ 5.3 $\mathtt{GradWhitening}$: An Efficient Non-diagonal Second-Order Update
Approximating Hessian with a Kronecker factorization is not new and has been extensively studied in the literature (Martens & Grosse, 2015; George et al., 2018; Gao et al., 2023; 2021; Koroko et al., 2022; Eschenhagen et al., 2024; Gupta et al., 2018). Here, this specific structure is useful in our context as 1) it enables estimation of Hessian/FIM without temporal EMAs, nor mini-batch statistics, and 2) it aligns with the statistical property of STB, as shown later in Proposition 1.
By leveraging assumption 1, we can now effectively estimate ${\mathbf{H}}$ by only using one single gradient matrix sample $\bm{G}:=[{\bm{g}}_{1},\dots,{\bm{g}}_{n}]\in\mathbb{R}^{m\times n}$. Recall that the Fisher information formulation of Hessian is defined as ${\mathbf{H}}=\mathbb{E}[\text{Vec}({\mathbf{G}})\text{Vec}({\mathbf{G}})^{\top}]$ where $\text{Vec}(\cdot)$ denotes the vectorized operation. Under assumption 1, we can estimate ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$ by computing the following simple estimate $\frac{1}{n}\sum_{i=1}^{n}{\bm{g}}_{i}{\bm{g}}_{i}^{\top}=\bm{G}\bm{G}^{\top}$, which approximate the current $\tilde{{\mathbf{H}}}$. Hence, $\mathtt{GradWhitening}$ can be seen as applying a second order update under our structural assumption:
@@
In the following Proposition, we show that the assumption 1 of $\mathtt{GradWhitening}$ aligns with the equilibrium Hessian structure in the STB regime.
@@
Proposition 1 shows that, under a simplified setting of the transformer, the Hessian will also converge to an equilibrium solution where the $M_{C}\times M_{C}$ blocks over the diagonal direction of Hessian shares an identical structure, which supports the assumption of $\mathtt{GradWhitening}$. This result is verified in our numerical experiment (Appendix, Figure 9). Finally, Theorem 2 presented in Appendix will also show how $\mathtt{GradWhitening}$ helps to make the convergence rate of SGD more robust to the condition number of local curvatures, and outperforms both SGD and Adam in the ill-conditioned regime.
@@ 6 Experiments
In this section, we report empirical results for SWAN. The main purposes are:
  • Benchmarking the performance, training speed-ups and memory footprint of SWAN on pre-training LLaMA with C4 dataset (Raffel et al., 2020) against Adam. Results are presented in Section 6.1, Section 6.2, and Section 6.3.
  • Investigating the contribution of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ to the performance of SWAN through ablation studies, as outlined in Section 6.4.
  • Separately evaluating whether $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ achieve their intended purposes of stabilizing gradient noise and modeling/ counteracting local curvature, respectively. Detailed results are reported in Section 6.5.1 and Section 6.5.2.